SUBMILLIMETER SPECTROSCOPIC DIAGNOSTICS IN A SEMICONDUCTOR PROCESSING PLASMA

\underline{YASER H. HELAL}, CHRISTOPHER F. NEESE, JENNIFER A. HOLT, FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210; PAUL R. EWING, Applied Materials, Austin, TX 78724; PHILLIP J. STOUT, MICHAEL D. ARMACOST, Applied Materials, Sunnyvale, CA 94085.

Millimeter and submillimeter rotational spectroscopy was used to characterize and measure the abundances of compounds in a semiconductor processing plasma. Plasmas were generated using flow mixtures of Ar, C_{4}F_{8}, and O_{2} in a chamber with quartz windows for submillimeter wave transmission. Species of interest included the plasma products CF, CF_{2}, COF_{2}, and CO. Abundances as a function of flow mixtures and pressures as well as rf drive levels will be presented.